OpenStudy (anonymous):

Solve for the variable. -0.5(r – 2) = -r – 2 r = ______

OpenStudy (anonymous):

distribute the -0.5 to the terms inside the ( ) -0.5(r – 2) = -r – 2 -0.5r + 1 = -r - 2 +r +r ____ ____ 0.5r + 1 = -2 -1 -1 _______ _____ 0.5r = -3 ---- ---- 0.5 0.5 r = -6